OpenStudy (anonymous):

The portion of a student’s ballpoint pen that contains the ink is a cylinder with a diameter of 0.400 cm and a height of 11.5 cm. If the ink lasts 7 weeks, what is the volume of ink that the student uses each week?

OpenStudy (anonymous):

for the volume of a cylinder=pi*r^2*h here diameter=0.400 so radius=0.400/2=0.200 so V of cylinder=pi*(0.200)^2*11.5 solve it you will get=1.444 now this remian for 7 weeks...to get ink used for 1 week just divide it by 7 =1.444/7=0.207